# Witnessing mass-energy equivalence with trapped atom interferometers

### v2

@@ Abstract
We propose an experimental setup to probe the interplay between the quantum superposition principle and the gravitational time dilation arising from the mass-energy equivalence. It capitalizes on state-of-the-art atom interferometers that can keep atoms trapped in a superposition of heights in Earth’s gravitational field for exceedingly long times reaching minute-scale. Our proposal consists of adding two additional laser pulses to the existing setups that would set up a clock trapped at a superposition of heights reading a quantum superposition of relativistic proper times. We develop a method to include relativistic corrections to Bloch oscillations which describe the trapped part of the interferometer. As a result, we derive all the trajectories arising in this setup, the phases acquired by different trajectories in the interferometer, and demonstrate that the effect of superposition of proper times would manifest itself in the interference pattern in two distinct ways: as visibility modulations, and as a shift of the resonant frequency of the atom. We argue that the latter might be observable with current technology.
@@
—Introduction. A key prediction of general relativity is that time is not a global background parameter but flows at different rates depending on the spacetime geometry, a phenomenon known as time dilation [1]. The physical effects produced by time dilation have been tested in numerous paradigmatic experiments. Pound and Rebka were the first to observe an altitude-dependent gravitational redshift in gamma rays emitted between the top and bottom of a tower [2]; Hafele and Keating directly tested special and general relativistic time dilation by comparing actual clocks at different heights and moving at different speeds [3]; while Shapiro proposed to measure the reduction of the speed of light for electromagnetic waves traveling across regions subject to a gravitational potential [4], which was observed in [5]. The above effects are fully explainable within the paradigm of classical (non-quantum) relativistic physics.
@@
The first experiment measuring the effect of gravity on the quantum wavefunction of a single particle was performed by Colella, Overhauser, and Werner (COW) [6]. The setup considered neutrons traveling in a superposition of heights in the Earth’s gravitational field, following the basic geometry of a Mach-Zehnder-type interferometer. The neutrons, traversing arms at differnt heights in the gravitational potential acquire a relative phase. The resulting interference pattern was measured and a phase difference consistent with the Earth’s gravitational potential was deduced. These and related effects measured with atoms [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22] are thus tests of the non-relativistic gravitational effects in quantum systems.
@@
Hence, none of the experiments performed so far probe the interplay between quantum mechanics and general relativity — the idea of how to do it was only developed in [23, 24]. These works considered interferometric experiments with a particle having an internal degree of freedom that could be treated operationally as a “quantum clock”. The internal evolution would then depend on the proper time along the paths taken through general curved spacetime [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34]. Specifically, the authors of [24] predicted that in the proposed experiment one would observe oscillations of the visibility of the interference pattern, defined as the contrast between the constructive and destructive interference fringes (which in the non-relativistic theory in the COW-type experiments is, in principle, always maximal). The predicted periodic modulations of the visibility would be an effect that arises unambiguously from the interplay between relativistic and quantum physics. Several proposals were made to probe this effect in interferometric experiments with electrons [35] and atoms [27, 28, 36, 37, 38]. However, given the size of the interferometer required for a reasonable visibility drop [24] it is rather unlikely that this effect will be observed anytime soon.
@@
Derivations of the dynamics of relativistic quantum particles with quantized internal mass-energy, necessary to make predictions for the above experiments, have been done in the framework of relativistic quantum mechanics in refs [39, 40, 41, 42], and starting from quantum field theory in curved background in refs [43, 44, 28]. In the low energy limit for the center of mass of the composite particles, they coincide. These results were applied to modeling the dynamics of particles in clock interferometers [35, 27, 28, 36, 45, 37] and of trapped particles [46], in particular in the context of optical clocks to explore relativistic contributions to the clock’s frequency and precision in refs [47, 48, 49], and most recently to explore associated many-body effects [50].
@@
In this work, we propose a concrete experimental setup for witnessing time dilation effects in atom interference. Our idea is inspired by the rapid development of trapped atom interferometers [18, 19, 20, 21, 22] utilizing Bloch oscillations to keep atoms in a coherent superposition of heights in a gravitational field for increasingly long times, reaching minute-scale [22]. We propose a modification of the existing interferometers in the form of two additional “clock pulses” at the beginning and the end of the Bloch oscillations. This procedure effectively sets up a clock at a superposition of heights in the gravitational field, which means that the clock experiences a superposition of proper times. We develop a method to account for relativistic mass-energy corrections in Bloch oscillations and use it to derive all the phases acquired by different trajectories within the interferometer and analyze the interference pattern. We recover the visibility modulations predicted in [24], where fixed classical trajectories of the clocks within the interferometer were assumed, but crucially, we predict the clock’s frequency shift, making this proposal feasible.
@@
Indeed, the proposed setup implements Ramsey spectroscopy for spatially superposed atoms, allowing for measurements of their resonant frequencies. We show that relativistic effects manifest therein as resonant-frequency shifts. We argue that this will be easier to observe than the visibility oscillations due to the enormous precision of frequency measurements [51, 52]. In particular, for the height separations reported in ref. [20], the predicted fractional frequency shift is of the order of $10^{-22}$, one order of magnitude below the current measurement precision. Revealing effects potentially detectable with current technology, our proposal constitutes a major step toward experimental tests of the interplay between quantum mechanics and general relativity.
@@
—Setup. Let us consider a one-dimensional atom interferometry setup, Fig. 1, closely resembling configurations implemented in [18, 19, 20, 21, 22]. It involves a highly cooled atom 111in experiments [20, 21, 22] the temperature reaches a few hundreds of nanokelvins moving along the gravitational field $g$ with velocity modified by Bragg pulses and an optical lattice. We assume a two-level internal structure of the atom with the ground and excited states separated by energy $\hbar\omega_{0}$.
@@
Figure 1: Proposed setup. All motion is restricted to one dimension (parallel to the gravitational field). Blue wavy lines represent $\pi/2$ Bragg pulses splitting the trajectory of the atom, while red wavy lines represent the $\pi/2$ clock pulses affecting only the internal state of the atom. The optical lattice (green striped pattern) is turned on exactly when the blue trajectories reach their apex and off after time $T_{\text{B}}$. Shaded dashed lines represent trajectories lost within the interferometer.
First, atoms are prepared in the ground state and launched upwards into free fall with initial velocity $v_{0}$. Next, a pair of $\pi/2$ Bragg pulses are applied separated by a time $T$. Each of them splits the atomic trajectory into two — one with unchanged velocity, and one with velocity modified by $\Delta v$. Therefore, after this sequence, the initial trajectory of each atom is split into four paths with pairwise equal velocities — one pair with velocity $v_{0}-gT$, and the other pair with velocity greater by $\Delta v$. We focus on the latter pair and assume that the track of the remaining trajectories is lost as in experiments in refs [18, 19, 20, 21, 22].
When the relevant trajectories reach the apex, namely at time $T+T^{\prime}=(v_{0}+\Delta v)/g$, the optical lattice is switched on, and atoms performs Bloch oscillations. At switching on the lattice we introduce our modification — a clock pulse at frequency $\omega$ that puts atoms in an equal superposition of the two internal states without affecting its momentum [54, 55, 56]. The optical lattice remains on for time $T_{\text{B}}$, after which the second clock pulse is applied and the lattice is switched off. The atoms fall freely for time $T^{\prime}$, after which another pair of $\pi/2$ Bragg pulses is applied to recombine the trajectory. At this stage, another two trajectories are lost — this is different than in experiments [18, 19, 20, 21, 22] where all the output ports are measured but only two interfere (a modification that we introduce for simplicity of the subsequent analysis).
In the end, one measures the probabilities $P_{g}^{(j)}$ that the atoms are found in the ground state in the output trajectory $j\in\{0,1\}$, with 0 and 1 corresponding to the red and blue trajectories in Fig. 1, respectively). Apart from the non-relativistic phase $\Delta\phi$ between the upper and lower ground state trajectories (measured in typical interferometric experiments, e.g. [20]), there will be additional phases $\delta_{\text{d}}$ and $\delta_{\text{u}}$ acquired within the optical lattice by the excited state on the lower and upper trajectory, respectively, which can be measured with the proposed setup. The probability $P_{g}^{(j)}$ is given by (see Appendix A for derivation)
$\begin{split}P_{g}^{(j)}=\frac{1}{16}&\bigg{[}1-\cos\left(\frac{\delta_{\text{u}}+\delta_{\text{d}}}{2}\right)\cos\left(\frac{\delta_{\text{u}}-\delta_{\text{d}}}{2}\right)+(-1)^{j}\times\\ &\times 2\cos\left(\Delta\phi+\frac{\delta_{\text{u}}-\delta_{\text{d}}}{2}\right)\sin\left(\frac{\delta_{\text{d}}}{2}\right)\sin\left(\frac{\delta_{\text{u}}}{2}\right)\bigg{]}.\end{split}$ (1)
One can get rid of the $\Delta\phi$-dependence by calculating the total probability $P_{g}=P_{g}^{(0)}+P_{g}^{(1)}$ of finding the atom in the ground state
$P_{g}=\frac{1}{8}\left[1-\cos\left(\frac{\delta_{\text{u}}-\delta_{\text{d}}}{2}\right)\cos\left(\frac{\delta_{\text{u}}+\delta_{\text{d}}}{2}\right)\right].$ (2)
This probability does not depend on the specific opening and closure of the interferometer, but only on the part confined in the optical lattice. This part is central to our considerations because the atom will effectively behave as a clock at a superposition of heights within the optical lattice. The presence of two clock pulses is the main modification to previous experimental implementations [18, 19, 20, 21, 22]. Let us notice, that without those pulses, the probability of measuring the atoms in the ground state would be constant and equal to 1/4 (recall that we lose some of the trajectories). In our setup, they oscillate between 0 and 1/4, depending on the time $T_{\text{B}}$ of the lattice hold.
—Dynamics and phases. Let us analyze the optical lattice stage of the interferometer and the additional phases $\delta_{\text{d}}$ and $\delta_{\text{u}}$ acquired by the excited state. At this stage, the atom will Bloch oscillate at a superposition of heights separated by $\Delta z$. We adopt a simplified model of Bloch oscillations (see [57, 19] and Fig. 2), in which the atom interacts with the optical lattice only when it reaches a downward velocity $v_{\text{B}}=\hbar k/m$ (here $k$ is the wave vector of the optical-lattice laser and $m$ is the mass of the ground-state atom). Then, it instantaneously exchanges two photons with the lattice and changes its velocity. For the appropriately chosen frequency of the lattice light, the velocity at the interaction point will simply flip its sign, and the atom will oscillate at the fixed height with period $\tau_{\text{B}}=2v_{\text{B}}/g$. We will assume that the laser frequency has been chosen exactly this way. Each bounce starts and ends at the apex of a free-fall trajectory. Within the time $T_{\text{B}}$ the atom performs $N$ oscillations with the last one possibly extended (or shortened) by a time $\tau\in(-\tau_{\text{B}}/2,\tau_{\text{B}}/2)$ (this is because $T_{\text{B}}$ is not necessarily an integer multiple of the Bloch period $\tau_{\text{B}}$; see Fig. 2).
$N\tau_{\text{B}}$$\tau$$T_{\text{B}}$
@@
Figure 2: Trajectory followed by the atom (at the lower or upper trajectory) in the optical lattice. We adopt a simplified model of Bloch oscillations in which the atom follows a bouncing trajectory at a fixed height. The total time of Bloch oscillations $T_{\text{B}}$ needs not be equal to an integer multiple of Bloch periods $N\tau_{\text{B}}$. Instead, $T_{\text{B}}=N\tau_{\text{B}}+\tau$ with $\tau\in(-\pi/2,\pi/2)$.
@@
Crucially, due to the mass difference between the excited and the ground state of the atom, the condition for bouncing at the fixed height can be satisfied only for one of the internal states for a given lattice. For instance, if we choose the lattice laser frequency so that the condition is satisfied for the ground state, the excited state will necessarily fall beyond the starting height with each oscillation. However, it can be shown (see Appendix B) that in viable experimental scenarios, the total fall is much smaller than the thermal wavelength of the atom. Therefore, employing the perturbative approach developed in [58] is justified, which here means that we can approximate that both the ground and the excited state follow the same trajectory. We further verify that this approximation holds in Appendix D.
@@
The phases $\delta_{\text{d}}$ and $\delta_{\text{u}}$ consist of two components: laser phase $\omega T_{\text{B}}$ acquired by the excited state due to the clock pulses, and a propagation phase correction due to different masses of the ground and the excited state. To calculate the latter we follow [58] and express the Lagrangian of the excited state $L_{e}=L_{g}+\delta L$ in terms of the Lagrangian $L_{g}$ of the ground state and a small correction $\delta L$ resulting from the rest-mass difference between the ground and the excited state. If we denote this mass difference by $\delta m=\hbar\omega_{0}/c^{2}$, the Lagrangian correction is given by
@@
Then, to calculate the propagation phase correction we integrate $\delta L$ over the trajectory followed by the atom and divide by $\hbar$. The full calculation is presented in the Appendix C (followed by a non-perturbative calculation in Appendix D). It leads to the following formulae for the phases $\delta_{\text{d}}$ and $\delta_{\text{u}}$:
@@
$\begin{split}&\delta_{\text{d}}=\left[\omega-\omega_{0}\left(1+\frac{\langle v^{2}\rangle}{2c^{2}}\right)\right]T_{\text{B}},\\ &\delta_{\text{u}}=\left[\omega-\omega_{0}\left(1+\frac{\langle v^{2}\rangle}{2c^{2}}+\frac{g\Delta z}{c^{2}}\right)\right]T_{\text{B}},\end{split}$ (4)
@@
where $\langle v^{2}\rangle=v_{\text{B}}^{2}/3$ is the mean square speed within a single oscillation.
@@
In principle, the mass-energy corrections to the interferometric phases are not the only corrections of the order of $1/c^{2}$ that are present in the proposed experiment — following [59] we could also include corrections coming from the higher-order expansion of the free-fall Lagrangian of the relativistic particle, and $1/c^{2}$ corrections to the laser phases coming from elastic interactions (Bragg and Bloch pulses). These depend only on the trajectory followed by the atom and thus, in the picture in which ground and excited states follow the same trajectories, affect only the interferometric phase $\Delta\phi$ leaving $\delta_{\text{k}}$ and $\delta_{\text{g}}$ unmodified. However, the $1/c^{2}$ corrections to $\Delta\phi$ are not the main focus of our considerations and we will not deal with them.
@@
On the contrary, any corrections to the clock-pulse laser phases would affect $\delta_{\text{k}}$ and $\delta_{\text{g}}$ if they were present. Let us notice, however, that the corrections to the laser phases calculated in [59] resulted from the assumption that the atom-light interactions are always resonant and the atom interacts with photons with laboratory-frame frequencies depending on its position and velocity. For our present purposes, we assume instead that the laser frequency of the clock pulses is fixed and the non-resonant processes are possible. Then, there are no corrections to the clock-pulse laser phases because the atom interacts with photons with the same laboratory-frame frequency regardless of its height and velocity.
@@
—Results and discussion. Let us denote $\varepsilon_{\text{k}}=\langle v^{2}\rangle/2c^{2}$ for the kinetic correction in Eq. (4) and $\varepsilon_{\text{g}}=g\Delta z/c^{2}$ for the gravitational correction, and denote by $\langle\omega_{0}\rangle=\omega_{0}\left(1+\varepsilon_{\text{k}}+\varepsilon_{\text{g}}/2\right)$ the (quantum mechanical) mean of the atomic internal frequency. Then, the probability of detecting the atom in the ground state (compare with Eq. (2)) can be written as
$P_{g}=\frac{1}{8}\left[1-\mathcal{V}(T_{\text{B}})\cos\left(\left[\omega-\langle\omega_{0}\rangle\right]T_{\text{B}}\right)\right],$ (5)
@@
where $\mathcal{V}(T_{\text{B}}):=\cos\left(\varepsilon_{\text{g}}\omega_{0}T_{\text{B}}/2\right)$ is the interferometric visibility. In a generic situation (away from the resonance), the visibility varies very slowly with time $T_{\text{B}}$ compared to $\cos\left(\left[\omega-\langle\omega_{0}\rangle\right]T_{\text{B}}\right)$. Therefore, Eq. (5) describes an interference pattern with a slowly varying envelope. Apart from the visibility modulations, Eq. (5) looks like a result of a Ramsey interferometry experiment for an atom with internal transition frequency $\langle\omega_{0}\rangle$. For fixed time $T_{\text{B}}$ one can scan through different clock pulse frequencies $\omega$ to find this resonant frequency. Thus, the relativistic effects coming from mass-energy equivalence can be measured in two ways: by observing the visibility oscillations, and by measuring the fractional frequency shift
@@
$\left\langle\frac{\delta\omega_{0}}{\omega_{0}}\right\rangle\equiv\frac{\langle\omega_{0}\rangle-\omega_{0}}{\omega_{0}}=\varepsilon_{\text{k}}+\frac{\varepsilon_{\text{g}}}{2}.$ (6)

### v3

@@ Abstract
We propose an experimental setup to probe the interplay between the quantum superposition principle and gravitational time dilation arising from the mass-energy equivalence. It capitalizes on state-of-the-art atom interferometers that can keep atoms trapped in a superposition of heights in Earth’s gravitational field for exceedingly long times, reaching the minute scale. Our proposal consists of adding two additional laser pulses to the existing experiments that would set up a clock trapped at a superposition of heights, reading a quantum superposition of relativistic proper times. We develop a method to include relativistic corrections to Bloch oscillations, which describe the trapped part of the interferometer. We derive the trajectories and corresponding phases acquired in each arm of the interferometer. We then show that a superposition of proper times manifests in the interference pattern in two ways: visibility modulations and a shift of the atom’s resonant frequency. We argue that the latter might be observable with current technology.
@@
—Introduction. A key prediction of general relativity is that time is not a global background parameter but flows at different rates depending on the spacetime geometry, a phenomenon known as time dilation [1]. The physical effects produced by time dilation have been tested in numerous paradigmatic experiments: Pound and Rebka were the first to observe an altitude-dependent gravitational redshift in gamma rays emitted between the top and bottom of a tower [2]; Hafele and Keating directly tested special and general relativistic time dilation by comparing atomic clocks at different heights and moving at different speeds [3]; and Shapiro proposed to measure the reduction in the speed of light for electromagnetic waves traveling across regions subject to a gravitational potential [4], an effect later observed in [5]. These effects are fully explainable in classical relativistic physics.
@@
The first experiment to measure the effect of gravity on the quantum wavefunction of a single particle was performed by Colella, Overhauser, and Werner (COW) [6]. In this setup, neutrons travel in a superposition of heights in Earth’s gravitational field, following the geometry of a Mach-Zehnder-type interferometer. The neutrons, traversing arms at different heights in the gravitational potential, acquire a relative phase. The resulting interference pattern was measured, yielding a phase shift consistent with Earth’s gravitational potential. These experiments, and related measurements with atoms [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22], thus test non-relativistic gravitational effects in quantum systems.
@@
The proposal to unambiguously probe the interplay between quantum mechanics and general relativity was developed in [23, 24]. These works considered interferometric experiments with particles having an internal degree of freedom that can be treated operationally as a “quantum clock”. The internal evolution would then depend on the proper time along the paths taken through curved spacetime [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34]. Specifically, the authors of [24] predicted that the experiment would exhibit oscillations in the interference visibility — defined as the contrast between constructive and destructive fringes — which, in non-relativistic COW-type treatments, is always maximal in principle. The predicted periodic modulations of the visibility would be an effect that arises unambiguously from the interplay between relativistic and quantum physics. Several proposals have been made to investigate this effect in interferometric experiments with electrons [35], and with atoms [27, 28, 36, 37, 38]. However, given the size of the interferometer required for a reasonable visibility drop [24], it is rather unlikely that this effect will be observed in the near future.
@@
Derivations of the dynamics of relativistic quantum particles with quantized internal mass-energy — necessary for predicting the above experiments — have been performed both within relativistic quantum mechanics [39, 40, 41, 42] and via quantum field theory on curved backgrounds [43, 44, 28]. In the low-energy limit for the center-of-mass motion of the composite particles, these two approaches coincide. These results have been applied to model particle dynamics in clock interferometers [35, 27, 28, 36, 45, 37] and in trapped-particle setups [46, 47]. In particular, they have been used in optical-clock studies to explore relativistic contributions to the clock’s frequency and precision in  [48, 49, 50], and, most recently, to explore associated many-body effects [51].
@@
In this work, we propose a concrete experimental setup for witnessing time dilation effects in atom interferometry. Our idea is inspired by the rapid development of trapped atom interferometers [18, 19, 20, 21, 22] utilizing Bloch oscillations to keep atoms in a coherent superposition of heights in a gravitational field for increasingly long times, reaching the minute scale [22]. We propose a modification of the existing interferometers by adding two “clock pulses” at the beginning and end of the Bloch oscillations, effectively placing a clock in a superposition of heights and thus of proper times. We develop a method to include relativistic mass–energy corrections in the Bloch oscillations and use it to derive all phases acquired along the interferometer trajectories. We then analyze the resulting interference pattern. We recover the visibility modulations predicted in [24], which assumed fixed classical trajectories of the clocks, and crucially predict a clock frequency shift that renders this proposal experimentally feasible.
@@
Indeed, the proposed setup implements Ramsey spectroscopy for spatially superposed atoms, allowing for measurements of their resonant frequencies. We show that relativistic effects manifest therein as resonant-frequency shifts. We argue that this will be easier to observe than the visibility oscillations thanks to the extreme precision of frequency measurements [52, 53]. For the height separations reported in ref. [20], the predicted fractional frequency shift is of order $10^{-22}$ — one order of magnitude smaller than current experimental precision. By revealing effects potentially detectable with existing technology, our proposal represents a major step toward experimental tests of the interplay between quantum mechanics and general relativity.
@@
—Setup. Let us consider a one-dimensional atom interferometry setup, Fig. 1, closely resembling configurations implemented in [18, 19, 20, 21, 22]. It involves highly cooled atoms [^1] In experiments [20, 21, 22], the temperature reaches a few hundred nanokelvins. moving along the gravitational field $g$, with their velocity manipulated by Bragg pulses and an optical lattice. We assume that atoms have a two-level internal structure, with ground and excited states separated by energy $\hbar\omega_{0}$.
$_{\rm B}TT$$^{\prime}TT$$T^{\prime}$
Figure 1: Proposed setup. All motion is restricted to one dimension (parallel to the gravitational field). Blue wavy lines represent $\pi/2$ Bragg pulses splitting the trajectory of the atom, while red wavy lines represent the $\pi/2$ clock pulses affecting only the internal state of the atom. The optical lattice (green striped pattern) is turned on exactly when the blue trajectories reach their apex and off after time $T_{\rm B}$. Shaded dashed lines represent trajectories lost within the interferometer.
Atoms are prepared in their ground state and launched upward with initial velocity $v_{0}$, entering free fall under gravity. Then, two $\pi/2$ Bragg pulses are applied, separated by an interval $T$. Each pulse splits the atomic trajectory into two: one retains its velocity, while the other is shifted by $\Delta v$. Consequently, each initial trajectory splits into four paths with two distinct velocities: two with $v_{0}-gT$ and two with $v_{0}-gT+\Delta v$. We focus on the latter two paths, assuming that the others are lost, as in [18, 19, 20, 21, 22].
When the relevant trajectories reach their apex at time $T+T^{\prime}=(v_{0}+\Delta v)/g$, the optical lattice is switched on, and atoms perform Bloch oscillations. When the lattice is switched on, we introduce our modification—a clock pulse at frequency $\omega$ that places atoms into an equal superposition of the two internal states without affecting their momentum [54, 55, 56, 57, 58, 59, 60]. The optical lattice remains on for a duration $T_{\mathrm{B}}$, after which the second clock pulse is applied and the lattice is turned off. The atoms then fall freely for time $T^{\prime}$, after which another pair of $\pi/2$ Bragg pulses (separated by an interval $T$) is applied to recombine the trajectories. At this stage, two additional trajectories are lost, unlike in [18, 19, 20, 21, 22], where all output ports are measured, although only two interfere (a simplification introduced here for clarity in the subsequent analysis).
In the end, one measures the probabilities $P_{g}^{(j)}$ that the atoms are found in the ground state in output trajectory $j\in\{0,1\}$ (with 0 and 1 corresponding to the red and blue trajectories in Fig. 1, respectively). Apart from the non-relativistic phase $\Delta\phi$ between the upper and lower ground-state trajectories (measured in typical interferometric experiments, e.g. [20]), there will be additional phases $\delta_{\rm d}$ and $\delta_{\rm u}$ acquired within the optical lattice by the excited state on the lower and upper trajectories and which can be measured with the proposed setup. The probability $P_{g}^{(j)}$ is given by (see Appendix A for derivation):
$P_{g}^{(j)}=\frac{1}{16}\bigg{[}1-\cos\left(\frac{\delta_{\rm u}+\delta_{\rm d}}{2}\right)\cos\left(\frac{\delta_{\rm u}-\delta_{\rm d}}{2}\right)+2(-1)^{j}\cos\left(\Delta\phi+\frac{\delta_{\rm u}-\delta_{\rm d}}{2}\right)\sin\left(\frac{\delta_{\rm d}}{2}\right)\sin\left(\frac{\delta_{\rm u}}{2}\right)\bigg{]}.$ (1)
To eliminate the $\Delta\phi$-dependence, we calculate the total probability $P_{g}=P_{g}^{(0)}+P_{g}^{(1)}$ of finding the atom in the ground state
$P_{g}=\frac{1}{8}\left[1-\cos\left(\frac{\delta_{\rm u}-\delta_{\rm d}}{2}\right)\cos\left(\frac{\delta_{\rm u}+\delta_{\rm d}}{2}\right)\right].$ (2)
This probability depends only on the evolution confined to the optical-lattice region, where the atom effectively behaves as a clock in a superposition of heights. The presence of two clock pulses is the sole modification to previous implementations [18, 19, 20, 21, 22]. Note that without these pulses, the ground-state probability would remain fixed at 1/4 (recall that we lose some of the trajectories). In our setup, they oscillate between 0 and 1/4 as a function of the time $T_{\rm B}$ between the clock pulses.
—Dynamics and phases. We now analyze the optical-lattice stage of the interferometer and the additional phases $\delta_{\rm d}$ and $\delta_{\rm u}$ acquired by the excited state. During this stage, the atom undergoes Bloch oscillations in a superposition of two heights separated by $\Delta z$. We adopt a simplified model of Bloch oscillations (see [61, 19] and Fig. 2), in which the atom interacts with the optical lattice only when its downward velocity reaches $v_{\rm B}=\hbar k/m$ (here $k$ is the wave vector of the optical-lattice laser and $m$ is the mass of the ground-state atom). At that point, the atom instantaneously exchanges two photons with the lattice. For an appropriately chosen lattice-light frequency, this leads to velocity reversal occurring always at the same height, causing the atom to oscillate with period $\tau_{\rm B}=2v_{\rm B}/g$. We will assume that the laser frequency has been chosen exactly this way. Each oscillation thus begins and ends at the apex of a free-fall trajectory. Over the hold time $T_{\rm B}$, the atom completes $N$ full oscillations, with the final one possibly extended or shortened by an additional time $\tau\in(-\tau_{\rm B}/2,\tau_{\rm B}/2)$, reflecting that $T_{\rm B}$ need not be an integer multiple of $\tau_{\rm B}$.
@@
Figure 2: Trajectory followed by the atom (at the lower or upper trajectory) in the optical lattice. We adopt a simplified model of Bloch oscillations in which the atom follows a bouncing trajectory at a fixed height. The total time of Bloch oscillations $T_{\rm B}$ need not be equal to an integer multiple of Bloch periods $N\tau_{\rm B}$. Instead, $T_{\rm B}=N\tau_{\rm B}+\tau$ with $\tau\in(-\tau_{\rm B}/2,\tau_{\rm B}/2)$.
@@
Crucially, due to the mass difference between the excited and ground states of the atom, the condition for oscillations at the fixed height can be satisfied only for one internal state at a time for a given lattice frequency. If the lattice is tuned to satisfy this condition for the ground state, the excited state will fall below the starting height after each oscillation. However, as shown in Appendix B, in viable experimental scenarios, the cumulative fall distance is much smaller than the atom’s thermal de Broglie wavelength. Therefore, the perturbative approach of [62] is justified, allowing us to approximate that both internal states follow the same trajectory. This approximation is further validated in Appendix D.
@@
The phases $\delta_{\rm d}$ and $\delta_{\rm u}$ comprise two contributions: the laser phase $\omega T_{\rm B}$ acquired by the excited state due to the clock pulses, and a propagation-phase correction arising from the different masses of the ground and excited states. To calculate the latter, we follow [62] and write the excited-state Lagrangian as $L_{e}=L_{g}+\delta L$, where $L_{g}$ is the ground-state Lagrangian and $\delta L$ is a small correction resulting from the rest-mass difference between the two states. If we denote this mass difference by $\delta m=\hbar\omega_{0}/c^{2}$, the Lagrangian correction is given by:
@@
To calculate the propagation-phase correction, we integrate $\delta L$ along the atom’s trajectory and divide by $\hbar$. The full derivation is given in Appendix C; a non-perturbative calculation is presented in Appendix D. This yields the expressions for the phases $\delta_{\rm d}$ and $\delta_{\rm u}$:
@@
$\begin{split}&\delta_{\rm d}=\left[\omega-\omega_{0}\left(1-\frac{\langle v^{2}\rangle}{2c^{2}}\right)\right]T_{\rm B},\\ &\delta_{\rm u}=\left[\omega-\omega_{0}\left(1-\frac{\langle v^{2}\rangle}{2c^{2}}+\frac{g\Delta z}{c^{2}}\right)\right]T_{\rm B},\end{split}$ (4)
@@
where $\langle v^{2}\rangle=v_{\rm B}^{2}/3$ is the mean square speed within a single oscillation.
@@
In principle, mass–energy corrections to the interferometric phases are not the only $\mathcal{O}(1/c^{2})$ effects present in our proposed experiment. Following Werner et al. [63], one could also include higher-order corrections from the expansion of the free-fall Lagrangian of the relativistic particle, and $\mathcal{O}(1/c^{2})$ corrections to the laser phases arising from elastic interactions (Bragg and Bloch pulses). Because these contributions depend only on the atomic trajectory, and since we approximate that both internal states follow the same path, they modify only the interferometric phase $\Delta\phi$, leaving $\delta_{\rm d}$ and $\delta_{\rm u}$ unmodified. However, $\mathcal{O}(1/c^{2})$ corrections to $\Delta\phi$ lie beyond the main focus of this work and will be neglected.
@@
Conversely, any corrections to the clock-pulse phases would affect $\delta_{\rm d}$ and $\delta_{\rm u}$ if present. Note that the laser-phase corrections calculated in [63] arise from assuming resonant atom-light interactions, with the atom interacting with photons whose laboratory-frame frequencies depend on its position and velocity. Here, we instead assume the clock pulses operate at a fixed laboratory-frame frequency, permitting non-resonant processes. Under this assumption, there are no corrections to the clock-pulse phases, since the atom interacts with photons at the same frequency regardless of its height or velocity.
@@
—Results and discussion. Let us introduce $\varepsilon_{\rm k}=\langle v^{2}\rangle/2c^{2}$ for the kinetic correction in Eq. (4) and $\varepsilon_{\rm g}=g\Delta z/c^{2}$ for the gravitational correction. Denoting the quantum-mechanical mean internal frequency by $\langle\omega_{0}\rangle=\omega_{0}\left(1-\varepsilon_{\rm k}+\varepsilon_{\rm g}/2\right)$, the ground-state detection probability (cf. Eq. (2)) can be written as:
$P_{g}=\frac{1}{8}\left[1-\mathcal{V}(T_{\rm B})\cos\left(\left[\omega-\langle\omega_{0}\rangle\right]T_{\rm B}\right)\right],$ (5)
@@
where $\mathcal{V}(T_{\rm B}):=\cos\left(\varepsilon_{\rm g}\omega_{0}T_{\rm B}/2\right)$ is the interferometric visibility. In a generic operating regime away from resonance, the visibility envelope varies slowly with time $T_{\rm B}$ compared to the rapid oscillations of $\cos\left(\left[\omega-\langle\omega_{0}\rangle\right]T_{\rm B}\right)$. Thus, Eq. (5) yields an interference pattern comprising fast oscillations modulated by a slowly varying envelope. Apart from the visibility modulations, Eq. (5) looks like a result of a Ramsey interferometry experiment for an atom with internal transition frequency $\langle\omega_{0}\rangle$. For a fixed $T_{\rm B}$, scanning the clock-pulse frequency $\omega$ reveals the resonance at $\omega=\langle\omega_{0}\rangle$. Consequently, relativistic effects arising from mass–energy equivalence can be probed in two ways: by observing visibility oscillations, and by measuring the resulting fractional frequency shift
@@
$\left\langle\frac{\delta\omega_{0}}{\omega_{0}}\right\rangle\equiv\frac{\langle\omega_{0}\rangle-\omega_{0}}{\omega_{0}}=-\varepsilon_{\rm k}+\frac{\varepsilon_{\rm g}}{2}.$ (6)